4. A survey done by the Bermuda Tourism Board reported that the average cruise ship visitor to Bermuda on one of Cruise Burmuda’s cruise tours spends $185 for on-shore tickets, drinks, and the like (with a standard deviation of $15). Assume the distribution of visitor spending is uniform, and described by the probability density function. You can see the uniform distribution below… f(x) = 1/50 for x values between $160 and $210 0 everywhere else …where x =values for the random variable “cruise tourist spending”, and the probability density function: 1/50 Use the equation to answer the following questions within your Excel file. Show your calculations. (In Excel please) a. What percentage of cruise ship visitors spend between $170 and $180?
Answer #1
f(x) = 1/50 , a= 160 , b= 210
P(170<x<180)= 0.2 20%
Formula used= 1/50*(180-170)